the voltage applied across a given parallel-plate capacitor is double. How is the energy stored in the capacitor affected
Answer #1

energy = U = 1/2 C V^2

if voltage is doubled energy = U1 = 1/2 * C * ( 2 V )^2

                                                     U1 = 4 * 1/2 C V^2 = 4 U = Energy increases four times

                                                 U1 = 4 U = energy becomes four times of initial energu